This exercise is based on summary statistics rather than raw data. This information is typically all that is presented in published reports. You can perform inference procedures by hand from the summaries. Use the conservative Option 2 (degrees of freedom the smaller of

n1 − 1 and n2 − 1)

for two-sample t confidence intervals and P-values. You must trust that the authors understood the conditions for inference and verified that they apply. This isn't always true.

Equip male and female students with a small device that secretly records sound for a random 30 seconds during each 12.5-minute period over two days. Count the words each subject speaks during each recording period, and from this, estimate how many words per day each subject speaks. The published report includes a table summarizing six such studies. Here are two of the six.

Readers are supposed to understand that, for example, the 56 women in the first study had

x = 16,177

and

s = 7520.

It is commonly thought that women talk more than men. Does either of the two samples support this idea? Do the following for each study.


(b) Find the two sample t statistic. (Round your answers to three decimal places.)

Study 1:     t =
Study 2:     t =


(c) What degrees of freedom does Option 2 use to get a conservative P-value?

Study 1:     df =
Study 2:     df =
Answer #1


a(16177-165690 (752092156 + 91082/56) Aus 1 : t = (x,-22) (5,^2/01+ gen 2 luz) Study 1 : t = b) -0.2484 study 2 :- = (16496-1